. You are bullish on Google stock. The current price is $1,075/share and you
have $10,0010 of your own to invest. You then borrow an additional $10,000
from your broker at an interest rate of 5% per year and invest a total of
$20,000 in the stock. If Google does not pay a dividend, which of the
following is TRUE?

A.) If Google stock declines by 20% over the next year, the total return on your
investment is -45%.

B.) All statements are true.

C.) Margin enables investors to achieve higher returns, however, there is greater
downside risk due to the leverage and interest cost.

D.) If Google stock goes up by 20% over the next year, the total return on your
investment is 35%.

E.) If Google stock remains unchanged over the year, your return on investment
would be -5%.

2. You are to receive a cash flow of $1,750 three years from now. Once
received, you will invest the money and earn a 5.5% return. What is the value
of the investment in year 5?

A.)$2,345

B.)$1,929

C.)$1,948

D.) $2,287

E.) $2,055

3. The yield on a 10-year US Treasury bond is 2.5%. ABC Co. and XYZ Co.
each issue a 10-year bond. ABC’s bond has a yield of 3.5%; while XYZ’s
bond has a yield of 5%. Select the statement below that BEST describes this
situation.

A.) The credit spread of ABC's bond is 1%; and XYZ Co. should have greater
default risk than ABC Co.

B.) XYZ Co. should have greater maturity risk than ABC Co.

C.)The credit spread of ABC’s bond is 3.5%. XYZ Co. should have a higher credit rating than ABC Co.

D.) The credit spread of XYZ’s bond is 5%.
Answer #1

Given:   invest a total of $20,000 in the stock of Google.

Borrowed from broker = $10,000

Own investment should be $10000 (typing error in question 10,0010 is wrong should be $10,000)

Interest paid to broker = 5% *10,000 = $500 annually

option A.) If Google stock declines by 20% over the next year, the total return on your investment is -45%.

Value of investment if Google stock goes down by 20% = 20000*0.8 = 16000

Interest to be paid to broker = $500

Loss from Google stock = (+16000-500-20000) = 4500

My return on investment will be = -(4500)/10,000 = -45% (note that -4500 is divided by 10,000 because that is the investment I have made.The other 10,000 has to be returned to the broker. So profit and losses are always calculated based on my investment)

D.) If Google stock goes up by 20% over the next year, the total return on your investment is 35%.

Value of investment if Google stock goes up by 20% = 20000*1.2 = 24000

Interest to be paid to broker = $500

Profit from Google stock = (24,000-20,000-500) = 3500

My return on investment will be = (3500)/10,000 = 35%

E.) If Google stock remains unchanged over the year, your return on investment would be -5%.

Value of investment if Google stock remains unchanged = 20,000

Interest to be paid to broker = $500

Profit from Google stock = (20,000-20,000-500) = -500

My return on investment will be = (-500)/10,000 = -5%

C.) Margin enables investors to achieve higher returns, however, there is greater downside risk due to the leverage and interest cost. It is evident from option A and option D.

Thus correct option is B.) All statements are true.

2) Cash received three years from now = $1750

Interest rate = 5.5% per annum

Value of investment at year 5 = 1750*(1+5.5%)^2 = 1947.79 = 1948

Option C is correct.

3) A.) The credit spread of ABC's bond is 1% ( 3.5% -2.5 % US Treasury bond) ; and XYZ Co. should have greater default risk than ABC Co ( As the yield for ABC bond is lower than XYZ bond.For two bonds the one with the higher yield tends to be more riskier that the one with lower yield.n )